# A Bayesian Dirichlet Auto-Regressive Conditional Heteroskedasticity Model for Forecasting Currency Shares

### v3

@@ Abstract
In marketplace finance, the daily mix of billing currencies is compositional data that drives forecasting, reporting, and treasury risk. We study Airbnb’s 2017–2020 currency‑fee shares across four regions and present a Bayesian Dirichlet ARMA model with a Dirichlet ARCH volatility component, B‑DARMA‑DARCH. The model keeps predictions on the simplex, models mean dynamics with VARMA terms on the additive log‑ratio scale, and lets precision vary over time so volatility can spike during disruptions and settle as conditions normalize. We include weekly and yearly seasonality and a trend, and we evaluate against Dirichlet ARMA and transformed‑Gaussian VARMA across simulation designs with misreported observations and temporary regime shifts, then validate on a held‑out window and test in 2020 Q4. Across simulations and the application, B‑DARMA‑DARCH delivers the most accurate aggregate forecasts, better‑calibrated intervals, and weaker residual persistence at medium lags. Modeling precision as a dynamic process provides a practical, interpretable way to forecast proportions and quantify uncertainty when the noise itself moves.
@@ 1 Introduction
Each day’s currency proportions must lie in a simplex (they are nonnegative and sum to one), a constraint that complicates standard time-series methods. One classical approach is to use log-ratio transformations, as pioneered by aitchison1982statistical and extended in works such as Cargnoni1997BayesianFO, ravishanker2001compositional, mills2010forecasting, and snyder2017forecasting, among others. Transformed additive log-ratio VARMA (tVARMA) models have been fruitfully applied to compositional data (brunsdon1998time; mills2009forecasting; mills2010forecasting).
@@
In contrast, the Dirichlet distribution models compositional data directly on the simplex. benjamin2003generalized proposed frequentist Generalized ARMA models that could be extended to the simplex. Contributions to Dirichlet time series include Bayesian compositional state space models by grunwald1993time, da2011dynamic, and da2015bayesian; the Bayesian Dirichlet Auto-Regressive Moving Average (B-DARMA) model by KATZ20241556; forecast7030032; forecast7040062; katz2025centeredmadirichletarma; and frequentist Dirichlet ARMA models by zheng2017dirichlet; creus2021dirichlet.
The generalized auto-regressive conditional heteroskedasticity (GARCH) model (bollerslev1986generalized) is fundamental in modeling financial time series volatility. Extensions include non-normal likelihoods to capture heavy tails and skewness (nelson1991conditional; engle2001theoretical). bauwens2006multivariate surveyed advancements in multivariate GARCH models.
@@ Acknowledgement
The authors thank Sean Wilson, Jess Needleman, Liz Medina, Yuling Kuo, Erica Savage, and Jackson Wang for helpful discussions, Adam Liss and Ellie Mertz for championing the research.
@@
The primary dataset used in our study, A Bayesian Dirichlet Auto-Regressive Conditional Heteroskedasticity Model for Forecasting Currency Shares, is not publicly available due to confidentiality constraints. However, the Stan code for the Bayesian Dirichlet Auto-Regressive Moving Average Dirichlet Auto-Regressive Conditional Heteroskedasticity Model (B-DARMA-DARCH) model is available for public access. It can be found at our GitHub repository: GitHub repository.
6 Supplementary Materials
Alternative Log-Ratio Transformations
Although we have used the additive log-ratio (ALR) link in equations (4), (5), and (6), other transformations to and from the simplex may be employed. A common choice is the centered log-ratio (clr), defined as
${\textnormal{clr}}(\mathbf{y})\;=\;\Bigl(\,\ln\bigl(\tfrac{y_{1}}{g(\mathbf{y})}\bigr),\;\ln\bigl(\tfrac{y_{2}}{g(\mathbf{y})}\bigr),\;\dots,\;\ln\bigl(\tfrac{y_{J}}{g(\mathbf{y})}\bigr)\Bigr),$
where $g(\mathbf{y})=(\,y_{1}\,y_{2}\,\cdots\,y_{J})^{1/J}$ is the geometric mean of $\mathbf{y}$. The clr mapping places all $J$ components in a transformed $\mathbb{R}^{J}$ space but introduces the constraint that their sum is zero.
Another option is to use the Isometric Log-Ratio (ilr) transformation. For each index $j$, the ilr-transformed coordinate can be expressed as
$z_{j}=\sqrt{\frac{r_{j}}{r_{j}+1}}\,\ln\biggl(\frac{g_{j}(\mathbf{y)}}{\bigl(\prod_{i\in H_{j}}y_{i}\bigr)^{1/r_{j}}}\biggr),$
where $g_{j}(\mathbf{y})$ is defined as the geometric mean of a chosen subset $S_{j}\subseteq\mathbf{y}$, and $H_{j}$ is the complement of $S_{j}$ in $\mathbf{y}$. The value $r_{j}$ is the number of elements in $H_{j}$. The particular way in which $S_{j}$ and $H_{j}$ are selected depends on the specific application and how one wishes to interpret the ilr coordinates.
These mappings are mutually related by linear transformations (egozcue2003isometric) and each provides a bijection from the (J1)(J-1)-dimensional simplex onto a (J1)(J-1)-dimensional subset of $\mathbb{R}^{J}$ or $\mathbb{R}^{J-1}$. Therefore, if the same priors and model assumptions are consistently adapted to each log-ratio space, the resulting compositional models remain equivalent—differing only in how one embeds the data in a Euclidean domain.
Definition of the Partial Autocorrelation Function (PACF)
The PACF at lag $k$ measures the partial correlation between the SSR at time $t$ and the SSR at time tkt-k, controlling for the effects of intermediate lags $1$ to k1k-1. It is defined as the last coefficient $\phi_{kk}$ in the autoregressive model of order $k$
$\text{SSR}_{t}=\phi_{k1}\text{SSR}_{t-1}+\phi_{k2}\text{SSR}_{t-2}+\dots+\phi_{k,k-1}\text{SSR}_{t-(k-1)}+\phi_{kk}\text{SSR}_{t-k}+\varepsilon_{t}$ (8)
where $\varepsilon_{t}$ is the white noise error term at time $t$. The coefficient $\phi_{kk}$ represents the PACF at lag $k$ and can be calculated recursively using the Yule-Walker equations
$\phi_{kk}=\frac{r_{k}-\sum_{j=1}^{k-1}\phi_{k-1,j}r_{k-j}}{1-\sum_{j=1}^{k-1}\phi_{k-1,j}r_{j}}$
Here, $r_{k}$ is the autocorrelation at lag $k$, and $\phi_{k-1,j}$ are the PACF coefficients from the previous lag k1k-1.
By examining the PACF of the SSRs, we can identify the presence of residual autocorrelation at various lags, which indicates whether the model has adequately captured the temporal dependencies in the data.
Standardized Residuals
For the B-DARMA and B-DARCH models, which are based on the Dirichlet distribution as defined in Section 2, we compute standardized residuals using the standard deviations derived from the Dirichlet parameters.
At each time t, the residuals are calculated as $\text{residual}_{tj}=y_{tj}-\mu_{tj},\quad\text{for }j=1,\dots,J.$ The variance of each component $y_{tj}$ under the Dirichlet distribution, Dirichlet($\phi_{t}\boldsymbol{\mu}_{t}$), is $\operatorname{Var}[y_{tj}]=\frac{\mu_{tj}(1-\mu_{tj})}{\phi_{t}+1}$, with standard deviation $\sigma_{tj}=\sqrt{\frac{\mu_{tj}(1-\mu_{tj})}{\phi_{t}+1}}.$
The standardized residuals are then computed as $\text{standardized\_residual}_{tj}=\frac{y_{tj}-\mu_{tj}}{\sigma_{tj}}$ and the sum of squared standardized residuals at time t is $\text{SSR}_{t}=\sum_{j=1}^{J}\left(\text{standardized\_residual}_{tj}\right)^{2}.$
For the B-tVARMA and B-TVP-tVARMA models, residuals are computed in the additive log-ratio (alr) transformed space, $\text{residual}_{tj}={\textnormal{alr}}(\mathbf{y}_{tj})-{\textnormal{alr}}(\mathbf{\hat{y}}_{tj}),$ where $j=1,\dots,4$. These residuals are standardized using the inverse of the Cholesky factor $L_{\Sigma}$ of the estimated covariance matrix $\Sigma$$\text{standardized\_residual}_{t}=L\Sigma^{-1}\times\text{residual}_{t}$. The sum of squared standardized residuals at time t is then calculated as $\text{SSR}_{t}=\sum_{j=1}^{J-1}(\text{standardized\_residual}_{tj})^{2}.$
By examining the PACF of $\{\text{SSR}\}_{t=1}^{T}$ we assess the presence of autocorrelation in the residuals. A well-specified model should produce residuals with minimal autocorrelation, indicating that temporal dependencies have been adequately captured.
Parameter Estimates
Supplementary Figure S13 shows the posterior densities of the auto-regressive ($\alpha$) and moving average ($\tau$) coefficients for each region. In Region 1, the posterior of $\alpha$ is concentrated between 0.25 and 0.75, peaking around 0.45, indicating moderate volatility persistence; shocks have a noticeable but not prolonged effect on volatility. In Region 2, the $\alpha$ coefficient has posterior mean around 0.8, reflecting strong dependence of current volatility on past volatility, and the $\tau$ coefficient centers around 2.75-2.75, indicating that large deviations from expected currency proportions significantly increase / decrease future volatility—characteristic of volatility clustering. Region 3 shows $\alpha$ tightly concentrated between 0.85 and 0.90 and $\tau$ centered around 1.2-1.2 suggesting very strong volatility persistence. In contrast, Region 4 exhibits complex dynamics: $\alpha_{1}$ is centered around $0.15$, indicating small persistence from the first lag; $\alpha_{2}$ is skewed positive and centered around $0.35$, suggesting a stronger influence from the second lag; $\alpha_{3}$ peaks slightly above 0.5, indicating moderate persistence; and the $\tau$ parameter is concentrated near $0.02$, implying minimal influence of past innovations.
Overall, the higher $\alpha$ values in Regions 2 and 3 indicate that volatility in currency fee proportions is more persistent in these regions. The negative $\tau$ coefficients suggest that large deviations—such as those caused by abrupt travel restrictions, policy changes, and fluctuating foreign exchange (FX) rates lead to increased future volatility. This implies that Regions 2 and 3 are more susceptible to external shocks affecting travel and currency markets, resulting in prolonged volatility in currency fee proportions. In contrast, Region 1, with moderate volatility persistence, tends to return to stability more rapidly after shocks, possibly due a more stable currency environment or a larger share of domestic travel. Region 4’s dynamics, characterized by less emphasis on immediate past volatility and innovations, may be attributed to regional factors such as diversified travel sources, differing policy responses, or economic structures.
Figure S1: Average PACF of the sum of squared standardized residual values across all 50 simulations for studies 1 & 2. B-DARCH consistently exhibits lower PACF values beyond lag 1, indicating reduced residual autocorrelation and suggesting that its dynamic precision component better captures volatility shocks in these simulated compositional time series.
Figure S2: Average PACF of the sum of squared standardized residual values across all 50 simulations for studies 3 & 4. The B-DARCH model (red) again shows smaller autocorrelation at higher lags, whereas B-DARMA (green) and B-tVARMA (red) leave more persistent structures in the residuals. This result underlines the advantage of modeling time-varying volatility when faced with abrupt changes in compositional data.
Figure S3: Airbnb data analysis- Proportion of fees by currency for four regions: weekly seasonal behavior. AUD is the Australian dollar, BRL is the Brazillian Real, CAD is the Canadian Dollar,CHF is the Swiss Franc, CLP is the Chilean Peso,EUR is the European Euro, GBP is the Great British Pound, MXN is the Mexican Peso, NZD is the New Zealand Dollar, and USD is the US Dollar. Strong weekly cycles appear in many currencies, motivating the inclusion of weekly seasonal terms in our model’s mean structure.
Figure S4: Airbnb data analysis - proportion of fees by billing currency for four regions from Jan 1, 2020 to June 30, 2020. AUD is the Australian dollar, BRL is the Brazillian Real, CAD is the Canadian Dollar,CHF is the Swiss Franc, CLP is the Chilean Peso,EUR is the European Euro, GBP is the Great British Pound, MXN is the Mexican Peso, NZD is the New Zealand Dollar, and USD is the US Dollar. Distinct spikes and dips across multiple currencies (e.g., sharp increase in BRL or EUR) reflect significant disruptions.
Figure S5: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR means for four regions from Jan 1, 2020 to June 30, 2020. USD is the reference component.
Figure S6: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR means for four regions from Jan 1, 2017 to June 30, 2020. USD is the reference component. Apart from seasonal patterns, we see notable multi-year trends and cyclical behaviors (e.g., increasing share of EUR in Region 2, upward drift in AUD in Region 3), underlining the presence of both trend and seasonality in compositional dynamics.
Figure S7: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR variance for four regions from Jan 1, 2017 to June 30, 2020. USD is the reference component.The data display periods of high variance that sometimes persist for weeks, demonstrating “volatility clustering.” These episodes justify the GARCH-like specification in our B-DARCH model, which allows the precision parameter $\phi_{t}$ to track spikes in uncertainty over time.
Figure S8: Airbnb data analysis- correlation between ALR and lagged ALR values for each currency in the four regions. USD is the reference component. High positive autocorrelations in many currencies (e.g., EUR, GBP) highlight the strong temporal persistence. Negative cross-lag correlations for some pairs (e.g., EUR vs. BRL in Region 4) indicate that certain currencies move in opposite directions. Both effects reinforce the need for a multi-component time-series framework with auto- and cross-dependence.
Figure S9: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 1.
Figure S10: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 2.
Figure S11: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 3.
Figure S12: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 4.
Figure S13: Airbnb data analysis: Densities of DARCH model coefficients for all 4 regions. Region 1 was modeled with a DARCH(1,0), Regions 2-3 modeled with a DARCH(1,1), and Region 4 modeled with a DARCH(3,1). Each plot shows the estimated density of a specific $\alpha_{l}$ (ar_phi) or $\tau_{k}$ (ma_phi) parameter. Regions 2 and 3 exhibit high $\alpha$ indicating persistent volatility once spikes occur, whereas Region 4’s coefficients suggest moderate or delayed response

### v4

@@ Abstract
In marketplace finance, the daily mix of billing currencies is compositional data that drives forecasting, reporting, and treasury risk. We study Airbnb’s currency-fee shares across four regions and present a Bayesian Dirichlet ARMA model with a time-varying precision component. The model keeps predictions on the simplex, captures mean dynamics on the additive log-ratio scale, and lets volatility spike during disruptions and settle as conditions normalize. We evaluate against standard Dirichlet and transformed-Gaussian alternatives using simulations with misreported observations and temporary regime shifts, then validate on held-out data. Across all settings, our approach delivers more accurate forecasts, better-calibrated intervals, and weaker residual persistence. Modeling precision as a dynamic process provides a practical, interpretable way to forecast proportions and quantify uncertainty when the noise itself moves.
@@ 1 Introduction
Each day’s currency proportions must lie in a simplex (they are nonnegative and sum to one), a constraint that complicates standard time-series methods. One classical approach is to use log-ratio transformations, as pioneered by Aitchison (1982) and extended in works such as Cargnoni,Claudia et al. (1997), Ravishanker et al. (2001), Mills (2010), and Snyder et al. (2017), among others. Transformed additive log-ratio VARMA (tVARMA) models have been fruitfully applied to compositional data (Brunsdon and Smith, 1998; Mills, 2009, 2010).
@@
In contrast, the Dirichlet distribution models compositional data directly on the simplex. Benjamin et al. (2003) proposed frequentist Generalized ARMA models that could be extended to the simplex. Contributions to Dirichlet time series include Bayesian compositional state space models by Grunwald et al. (1993), da-Silva et al. (2011), and da-Silva and Rodrigues (2015); the Bayesian Dirichlet Auto-Regressive Moving Average (B-DARMA) model by Katz et al. (2024, 2025); Katz and Maierhofer (2025); Katz (2025); and frequentist Dirichlet ARMA models by Zheng and Chen (2017); Creus-Martí et al. (2021).
The generalized auto-regressive conditional heteroskedasticity (GARCH) model (Bollerslev, 1986) is fundamental in modeling financial time series volatility. Extensions include non-normal likelihoods to capture heavy tails and skewness (Nelson, 1991; Engle, 2001). Bauwens et al. (2006) surveyed advancements in multivariate GARCH models.
@@ Acknowledgement
The authors thank Sean Wilson, Jess Needleman, Liz Medina, Yuling Kuo, and Jackson Wang for helpful discussions, Adam Liss and Ellie Mertz for championing the research, and the Editor-in-Chief, Associate Editors, and two anonymous reviewers for constructive feedback that improved the manuscript.
@@
The primary dataset used in our study, A Bayesian Dirichlet Auto-Regressive Conditional Heteroskedasticity Model for Forecasting Currency Shares, is not publicly available due to confidentiality constraints. However, the Stan code for the Bayesian Dirichlet Auto-Regressive Moving Average Dirichlet Auto-Regressive Conditional Heteroskedasticity Model (B-DARMA-DARCH) model is available for public access as well as the R scripts for the simulation studies. It can be found at our GitHub repository: GitHub repository.